OpenStudy (anonymous):

5 men and 7 boys can do a piece of work in 24 days working 10 hours a day.3men can do as much as 5 boys .in how many days will 9 men and 8 boys working 10 hours a day finish the same work

OpenStudy (anonymous):

m-men,b-boys 5m+7b=1/240 3m=5b so 15b=1/240

OpenStudy (anonymous):

9men=15boys so 15+8 =23boys 23b=1/10x

OpenStudy (anonymous):

15*24/23=ans

OpenStudy (anonymous):

@ganeshie8 help plse

OpenStudy (anonymous):

hours are same you should only consider days

ganeshie8 (ganeshie8):

yep ! just get everything into `men` maybe

OpenStudy (anonymous):

work of 3men=5 boys work of 5 men=\[\frac{ 5 }{ 3}\times 5=\frac{ 25 }{ 3 } boys\] work of 5 men+7 boys=\[\frac{ 25 }{ 3 }+7=\frac{ 25+21 }{ 3 }=\frac{ 46 }{ 3 } boys\] again work of 9 men\[\frac{ 5 }{ 3 }\times 9=15 boys\] work of 9 men +8 boys=15+8=23 boys

OpenStudy (anonymous):

can you complete it?

OpenStudy (anonymous):

15*24/23 IS IT right? plse tell me

OpenStudy (anonymous):

\[\frac{ 46 }{ 3 }~boys~ working~ 10 ~hrs~complete ~work~\in~days=24 \] \[23~boys~working~10~hrs~will~complete~work=24\times \frac{ 46 }{ 3 }\times \frac{ 1 }{ 23 }=?\]